OpenStudy (anonymous):

how to solve (y-7) - (y+7) = 3y

OpenStudy (anonymous):

y-7-y-7=3y -14=3y can you continue ?

OpenStudy (anonymous):

Distributing the negative sign you get: y-7-y-7=3y which is -14=3y divide by 3 and you get y= -14/3

OpenStudy (anonymous):

thank you for helping